A heavy box of mass 20 kg is pulled on a horizontal surface.If the coefficient of kinetic friction between the box and the horizontal surface is 0.25 find the force of friction exerted by the horizontal forces on the box

Grade:12

2 Answers

Askiitians Expert Bharath-IITD
14 years ago

Dear Pallavi,

The kinetic friction generated due the horizontal forces on the box is given by the formula Fk= μkmg = 0.25 * 20 * 9.8 N = 49N

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best  !!!

 


Regards,

Askiitians Experts

Adapa Bharath

Shankararama Sharma
13 years ago

While in motion only the kinetic friction acts. f=uN.

                                                               f=umg

                                                                =0.25*20*9.8

                                                                =49N

so the friction exerted by the horizontal forces is 49N.